OpenStudy (anonymous):

The cost, C, of producing n clocks can be found using the formula 21) C = 36 n + 81. What is the cost when no clocks are produced?

OpenStudy (anonymous):

n=0 right? so its 81? or is it 0?

OpenStudy (anonymous):

Yeah. Plug in 0 for n.

OpenStudy (anonymous):

The answer is 324 but I don't understand how they got it. Its roots and radicals.

OpenStudy (anonymous):

The cost, C, of producing n clocks can be found using the formula C = 36 n + 81. What is the cost when no clocks are produced?

OpenStudy (anonymous):

Do they provide an explanation?

OpenStudy (anonymous):

it doesnt seem to make sense there

OpenStudy (anonymous):

No there is a radical sign in front of the n +81

OpenStudy (anonymous):

\[36\sqrt{n+81}\] ? is that it?

OpenStudy (anonymous):

Oh ok. That makes sense. \[C = 36\sqrt{n + 81}\]\[C = 36\sqrt{0 + 81}\]\[C = 36\sqrt{81}\]\[C = 36(9)\]\[C = 324\]

OpenStudy (anonymous):

here we go thats it!

OpenStudy (anonymous):

thanks Calcmathlete! now I get it!!

OpenStudy (anonymous):

n :)

OpenStudy (anonymous):

ahhh there it is haha all probs need all the pieces

OpenStudy (anonymous):

lol *np :)

OpenStudy (anonymous):

LOL!

OpenStudy (anonymous):

LQTMS!!!!!!!

OpenStudy (anonymous):

it the first button